# How can Diffusion Models Evolve into Continual Generators?

### v2

@@ Abstract
While diffusion models have achieved remarkable success in static data generation, their deployment in streaming or continual learning (CL) scenarios faces a major challenge: catastrophic forgetting (CF), where newly acquired generative capabilities overwrite previously learned ones. To systematically address this, we introduce a formal Continual Diffusion Generation (CDG) paradigm that characterizes and redefines CL in the context of generative diffusion models. Prior efforts often adapt heuristic strategies from continual classification tasks but lack alignment with the underlying diffusion process. In this work, we develop the first theoretical framework for CDG by analyzing cross-task dynamics in diffusion-based generative modeling. Our analysis reveals that the retention and stability of generative knowledge across tasks are governed by three key consistency criteria: inter-task knowledge consistency (IKC), unconditional knowledge consistency (UKC), and label knowledge consistency (LKC). Building on these insights, we propose Continual Consistency Diffusion (CCD), a principled framework that integrates these consistency objectives into training via hierarchical loss terms $\mathcal{L}_{IKC}$, $\mathcal{L}_{UKC}$, and $\mathcal{L}_{LKC}$. This promotes effective knowledge retention while enabling the assimilation of new generative capabilities. Extensive experiments on four benchmark datasets demonstrate that CCD achieves state-of-the-art performance under continual settings, with substantial gains in Mean Fidelity (MF) and Incremental Mean Fidelity (IMF), particularly in tasks with rich cross-task knowledge overlap.
1 Introduction
The remarkable success of diffusion models in synthesizing high-fidelity images has significantly accelerated the arrival of generative artificial intelligence (AGI). However, the static nature of their training paradigm hinders further advancement in dynamic real-world scenarios, such as personalized content creation platforms or real-time virtual environment generation for interactive applications. When faced with the arrival of new data, existing approaches generally involve retraining models on both historical and current data to prevent significant performance degradation or generative collapse  [^1] During the training of streaming tasks, we observe that diffusion models occasionally experience sudden catastrophic failures on certain tasks, leading to a significant degradation in performance. on previously learned tasks. However, this is computationally expensive and results in considerable waste of both computational resources and energy, exacerbating the practical challenges of deploying diffusion models in continual learning (CL) contexts. As a result, the need for an effective and stable Continual Diffusion Generation (CDG) framework has become increasingly urgent.
@@ 1 Introduction
Figure 1: An overview of Continual Diffusion Generation and its associated challenges. Current continual classification methods are largely ineffective, and often counterproductive, in preventing performance degradation and generative collapse in streaming tasks with diffusion models.
@@
While recent efforts have proposed a variety of promising strategies to mitigate catastrophic forgetting (CF) phenomenon in generative models, most remain insufficiently aligned with the core dynamics of diffusion-based architectures. In particular, many approaches adopt heuristics from continual classification tasks [33, 25, 34, 46, 20], or borrow stability-plasticity trade-off mechanisms originally developed for Generative Adversarial Networks (GANs) [1, 41, 42, 6, 48]. These techniques often conflict with the stochastic differential equations (SDEs) that govern the generative process in diffusion models, leading to noticeable degradation in output quality. Empirical results confirm these limitations, showing consistent declines in generative fidelity across sequential tasks reflected in metrics such as Fréchet Inception Distance (FID) (see Fig.1 and Tab.1). Therefore, to improve this situation, we begin by analyzing the mechanisms inherent in diffusion models and explore how they should be optimized in the context of streaming tasks to preserve shared knowledge across tasks as much as possible, thereby mitigating CF phenomenon in CDG.
@@
To achieve this goal, we begin by formalizing diffusion trajectories, complex sequences of denoising operations that progressively transform noise into structured outputs, and examining how they interact through shared knowledge under streaming task scenarios. Grounded in Bayesian theory and multi-task learning principles [44], we derive a theoretical upper bound that quantifies the retention of generative knowledge across tasks in CDG, as formally defined in Theorem 4. Attaining this upper bound requires satisfying three critical consistency constraints: inter-task alignment of model-internalized knowledge, consistency in the mean embeddings of unconditional samples across tasks, and semantic consistency within the prior (i.e., label) space of original samples  [^2] Notably, the label space in our setting serves as a proxy for human prior knowledge, and the constraint is not imposed to improve classification performance, but rather to preserve shared semantic structure across tasks. across tasks.
@@
To operationalize the theoretical insights, we introduce the Continual Consistency Diffusion (CCD) framework, which translates the derived guarantees into a tractable, consistency-driven optimization strategy. CCD enforces cross-task stability through a hierarchical integration of consistency objectives. Specifically, Inter-task Knowledge Consistency regularizes model-internal knowledge representations across tasks, serving as the foundation for long-term retention. Unconditional Knowledge Consistency preserves intrinsic generative behavior in the absence of explicit human priors, ensuring temporal coherence in the denoising process. Complementarily, Label Knowledge Consistency promotes alignment in the label space by enforcing semantic correspondence between original samples across tasks. Collectively, these mechanisms move beyond standard regularization or classification heuristics. By directly constraining the geometric structure of the diffusion process, CCD enables robust and theoretically grounded continual generation.
@@
In doing so, our work bridges the critical gap between traditional static diffusion models and the dynamic nature of real-world data streams. We present three main contributions. First, to the best of our knowledge, we are the first to rigorously formulate CDG from a diffusion perspective, rather than repurposing continual classification methods as done in previous studies. Second, we establish the first theoretical framework for CDG rooted in stochastic calculus, introducing novel stability bounds for SDE trajectories under sequential task adaptation (Theorem 4). Third, guided by these insights, we propose the CCD framework, which enforces intrinsic knowledge consistency through three synergistic components. Experiments on four benchmarks demonstrate its superiority, yielding significant gains while effectively mitigating CF phenomenon and generative collapse.
2 Related Works
@@
Diffusion Models. Diffusion-based generative architectures have redefined state-of-the-art performance in structured data synthesis, primarily due to their ability to invert stochastic denoising trajectories. At a foundational level, these models learn to reverse-engineer discrete Markov chains (DDPMs) [9] or continuous-time stochastic differential equations (SDEs) [32]. Recent advancements in noise scheduling [22, 19] and adaptive sampling techniques [19, 49] have further enhanced output fidelity. Such developments underscore diffusion models’ theoretical strength as universal data approximators [31], regardless of their discrete or continuous formulation. However, their success heavily relies on closed-world assumptions, where training data remains static and entirely observable. Consequently, a critical challenge persists in adapting diffusion models to dynamic, streaming data environments, paralleling incremental human cognition.
@@
Continual Learning. Continual Learning (CL) aims to enable models to progressively acquire new knowledge while retaining previously learned information, addressing the challenge of CF [12, 16, 24]. Traditional CL methodologies include replay-based techniques, which store subsets of historical data to maintain stable performance [26], and regularization methods like EWC [12] and LwF [16], which impose constraints on parameters to reduce interference between tasks. Additionally, gradient-based strategies such as GEM [18] orthogonalize gradients to minimize task conflicts. Recent advancements have increasingly emphasized memory efficiency [26] and robustness against domain shifts [24]. In parallel, the advent of pre-training has driven progress in parameter-efficient fine-tuning techniques. Methods such as L2P [40] and DualPrompt [39] utilize task-specific prompts to effectively balance adaptability and knowledge retention. Techniques like S-Prompt [38] and CODA-Prompt [28] enhance performance by explicitly capturing domain relationships, while dynamic methods like DAP [11] and hierarchical approaches like HiDe-Prompt [37] support adaptation across diverse domains. Innovations such as EASE [50] and NTK-CL [17] introduce expandable adapters and hybrid architectures to achieve efficient parameter tuning. Despite these advancements, the existing CL research remains focused on basic classification tasks, limiting applicability to complex real-world scenarios. To bridge this gap, we investigate extending CL methods to practical and challenging applications, particularly in the context of continual diffusion generation.
@@
Continual Diffusion Generation. CDG represents a significant blind spot in the current landscape of CL research, with only a scant body of work dedicated to this area. Among these studies, most have primarily explored methods based on GAN architectures [1, 41, 42, 6, 48]. Additionally, many approaches attempt to adapt techniques originally developed for continual classification problems [33, 25, 34, 46, 20], applying relevant fine-tuning strategies. However, given that generation and classification are fundamentally distinct tasks, such direct transfer is largely impractical. Our experimental findings further highlight that many methods effective in continual classification scenarios fail entirely in generation settings, sometimes even yielding adverse effects. This is primarily due to the substantial knowledge disparity across tasks, which causes the diffusion generator to collapse, a phenomenon we refer to as generative collapse. In this paper, we address this gap by developing a theoretical framework to model the task transition process of diffusion models.
Continual Diffusion Generation. CDG remains a significantly underexplored area within the broader field of CL, with only limited research dedicated specifically to this domain. Among the existing works, the majority have focused on GAN-based approaches [1, 41, 42, 6, 48], while others have attempted to adapt techniques originally designed for continual classification tasks [33, 25, 34, 46, 20] by applying fine-tuning strategies to generative settings. However, this cross-domain transfer is fraught with challenges. Generation and classification differ fundamentally in objective and structure, making many classification-based methods ill-suited for generative scenarios. Our empirical studies reinforce this observation, revealing that methods successful in continual classification often fail, or even harm performance, when directly applied to generative tasks. A central issue is the substantial knowledge shift across generative tasks, which destabilizes the learning dynamics of diffusion models and leads to what we term as generative collapse, a breakdown in the model’s ability to synthesize meaningful outputs. Recent advances have hinted at alternative pathways. For example, [43] shows that diffusion models could maintain high generative quality when trained on temporally correlated video streams, even with minimal continual mechanisms such as small replay buffers. This finding suggests that some degree of continuity in input distribution, such as temporal redundancy in video data, can naturally reinforce memory and help mitigate forgetting. Nonetheless, such strategies offer only coarse-grained control and lack the precision necessary to manage how generative knowledge is retained or adapted across tasks. This raises a key research question: how can we design principled mechanisms that provide fine-grained control over knowledge consolidation during task transitions in diffusion-based generative systems? In this paper, we tackle this challenge by introducing a theoretical framework that explicitly models task transitions in diffusion models. Our framework not only advances the theoretical understanding of CDG but also paves the way for developing more robust, adaptive, and scalable continual generative systems.
3 Continual Consistency Diffusion
@@
In CDG, we consider a sequence of non-stationary tasks $\{\mathcal{T}_{k}\}_{k=1}^{K}$, where each task $\mathcal{T}_{k}$ is associated with a distinct data distribution $p^{k}(x_{0})$ and a corresponding conditional label distribution $p^{k}(y|x_{0})$. The forward diffusion process for each task is governed by an SDE: $dx_{t}^{k}=f_{k}(x_{t}^{k},t)dt+g_{k}(t)dw_{t}$, where $x_{t}^{k}\in\mathbb{R}^{d}$ represents the diffused samples at time $t$ under task $\mathcal{T}_{k}$. To build on the derivations in Appendix B and enhance the preservation of shared knowledge across tasks, we employ a direct rehearsal buffer $\mathcal{B}^{real}_{k}=\{\hat{x}_{0},\hat{y}\sim p^{j}(x_{0},y)\}_{j=1}^{k-1}$,  [^3] We avoid generative replay as it requires generating past samples for each training batch, significantly extending optimization time by two to three times, making it impractical despite its performance benefits. which stores a limited set of real samples within a fixed storage budget $C$. This memory mechanism, both in its theoretical formulation and practical implementation, forms the core foundation for the retention of shared knowledge across tasks.
3.1 Theoretical Foundation
@@
A central challenge in CDG is formalizing the interaction of task-specific generative processes through shared SDE dynamics. Existing empirical approaches [30, 48, 46] mainly adapt solutions from continual classification methods, but lack rigorous theoretical guarantees, often leading to CF or rigid fixation. Our analysis begins by establishing fundamental bounds on cross-task knowledge retention, which are essential for systematic CDG. Through rigorous derivation, we obtain Theorem 4.
Theorem 1 (Cross-Task Diffusion Evolution Bound).
 [^4] For a rigorous proof, please refer to Appendix A and Appendix B.
@@
Let $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ represent two tasks in CDG, each characterized by distinct data distributions $p(x_{0})$ and $q(x_{0})$, along with their respective conditional label distributions $p(y)$ and $q(y)$. The forward processes for these tasks evolve over time as $\{p(x_{t})\}_{t=0}^{T}$ and $\{q(x_{t})\}_{t=0}^{T}$. Assume that for all $x_{0},y,t$, the conditional probability distributions of the two tasks satisfy $p_{t}(x_{t}|x_{0},y)=q_{t}(x_{t}|x_{0},y)$. Here, $p_{t}(x_{t}|x_{0},y)$ refers to the conditional distribution $p(x_{t}|x_{0},y,t)$, and similarly, $q_{t}(x_{t}|x_{0},y)$ denotes $q(x_{t}|x_{0},y,t)$. Furthermore, let $\epsilon_{\theta}^{p}$ and $\epsilon_{\theta}^{q}$ represent the time-dependent score approximators, or noise estimators, for tasks $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. Under mild assumptions, we expect that the gradients of the mean functions $\mu(x_{t},t)$ and $\nu(x_{t},t)$ align, such that $\nabla_{x_{t}}\mu(x_{t},t)\approx\nabla_{x_{t}}\nu(x_{t},t)$, as noted in [44]. Additionally, it is assumed that the variance $\sigma^{2}_{t}$ at any given time $t$ remains consistent across both tasks. Lastly, we assume that the evolving state $x_{t}$ does not influence the label $y$, meaning the label is independent of the diffusion process at any given time step. These conditions enable the potential retention and transfer of knowledge between tasks, leading to the derivation of an optimization upper bound for their interaction.
@@
$\mathcal{L}_{UB}\;=\;\kappa\mathcal{L}_{IKC}\;+\;\lambda\,\mathcal{L}_{UKC}\;+\;\eta\,\mathcal{L}_{LKC},$ (1)
@@
$\mathcal{L}_{IKC}\;=\;\epsilon_{\theta}^{q}\;-\;\epsilon_{\theta}^{p},$ (2)
@@
$\mathcal{L}_{LKC}\;=\;\frac{\bar{\alpha}_{t}}{\bar{\beta}_{t}}\mathbb{E}_{p_{t}(x_{0}|x_{t})}\Bigl{[}D_{\mathrm{KL}}\bigl{(}p_{t}(y|x_{0})\,\big{\|}\,q_{t}(y|x_{0})\bigr{)}\Bigr{]}.$ (4)
In particular, $\mathcal{L}_{UB}$ encapsulates three components: the inter-task knowledge consistency ($\mathcal{L}_{IKC}$), the unconditional knowledge consistency ($\mathcal{L}_{UKC}$), and the label knowledge consistency ($\mathcal{L}_{LKC}$). Minimizing $\mathcal{L}_{UB}$ aligns the reverse-time diffusion gradients between tasks $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$, thereby allowing the two tasks to retain as much shared knowledge as possible during the SDE optimization process.
Theorems 4 demonstrates that CF in CDG can be mitigated by enforcing consistency in the model’s internalization of knowledge, the mean of unconditional sample embeddings, and semantic information across original samples. To operationalize these principles, we develop a hierarchical optimization strategy that decomposes the theoretical bounds into tractable loss terms while preserving their information-theoretic foundations. First, we introduce the basic diffusion training loss and $\mathcal{L}_{IKC}$.
3.2 Basic Diffusion Model Training
To ensure the optimality of the score estimators $\epsilon_{\theta}^{p}$ and $\epsilon_{\theta}^{q}$ within $\mathcal{L}_{IKC}$ for effective subsequent retention and transfer, we define the fundamental training objective for each task $\mathcal{T}_{k}$. Following the standard DDPM framework [9, 32], the base objective for conditional generation minimizes the weighted $L_{2}$ error between predicted and actual noise:
@@ 3.2 Basic Diffusion Model Training
where $\bar{\alpha}_{t}$ and $\bar{\beta}_{t}$ follow the DDPM variance schedule [9], and $\epsilon$ represents standard Gaussian noise.
Building on this, to enhance the shared knowledge between different tasks, we incorporate the data from the buffer into the training process. However, instead of random sampling, we concatenate the pairs $<x_{0},\hat{x}_{0}>$, which also facilitates the implementation of Equations 2 to 4. Therefore, in the CDG, the fundamental composite training objective for diffusion models can be expressed as:

### v3

# CCD: Continual Consistency Diffusion for Lifelong Generative Modeling

@@ Abstract
While diffusion-based models have shown remarkable generative capabilities in static settings, their extension to continual learning (CL) scenarios remains fundamentally constrained by Generative Catastrophic Forgetting (GCF). We observe that even with a rehearsal buffer, new generative skills often overwrite previous ones, degrading performance on earlier tasks. Although some initial efforts have explored this space, most rely on heuristics borrowed from continual classification methods or use trained diffusion models as ad hoc replay generators, lacking a principled, unified solution to mitigating GCF and often conducting experiments under fragmented and inconsistent settings. To address this gap, we introduce the Continual Diffusion Generation (CDG), a structured pipeline that redefines how diffusion models are implemented under CL and enables systematic evaluation of GCF. Beyond the empirical pipeline, we propose the first theoretical foundation for CDG, grounded in a cross-task analysis of diffusion-specific generative dynamics. Our theoretical investigation identifies three fundamental consistency principles essential for preserving knowledge in the rehearsal buffer over time: inter-task knowledge consistency, unconditional knowledge consistency, and prior knowledge consistency. These criteria expose the latent mechanisms through which generative forgetting manifests across sequential tasks. Motivated by these insights, we further propose Continual Consistency Diffusion (CCD), a principled training framework that enforces these consistency objectives via hierarchical loss functions: $\mathcal{L}_{IKC}$, $\mathcal{L}_{UKC}$, and $\mathcal{L}_{PKC}$. This framework fosters long-term retention of generative knowledge and stable integration of new capabilities. Extensive experiments show that CCD achieves state-of-the-art performance across various benchmarks, especially improving generative metrics in overlapping-task scenarios.
Introduction
@@
Figure 1: An overview of CDG pipeline and its challenges, highlighting the ineffectiveness of current continual classification strategies in preventing performance degradation and GCF in streaming tasks with diffusion models.
The remarkable success of diffusion models in synthesizing high-fidelity text and images (Bruce et al. 2024; Nie et al. 2025; Yang et al. 2025) has significantly accelerated the arrival of generative artificial intelligence (AGI). However, the static nature of their training pipeline hinders further advancement in dynamic real-world scenarios, such as personalized content creation or real-time virtual environment generation for interactive applications. When faced with the arrival of new data, existing approaches generally involve retraining models on both historical and current data to prevent significant performance degradation or generative collapse  [^1] During the training of streaming tasks, we observe that diffusion models occasionally experience sudden catastrophic failures on certain tasks, leading to a significant degradation. on previously learned tasks. However, this is computationally expensive and results in considerable waste of both computational resources and energy, exacerbating the practical challenges of deploying diffusion models in continual learning (CL) contexts. These challenges are further compounded by the fact that current research on diffusion generation in CL remains fragmented, lacking a standardized experimental protocols (Zając et al. 2023; Masip et al. 2023; Cheng et al. 2024) and thorough understanding. This underscores the urgent need for a systematic pipeline capable of quantifying the corresponding continual generative performance in a standardized manner. Accordingly, we present a formally grounded Continual Diffusion Generation (CDG) pipeline (see Figure. 1), upon which our theoretical analysis and experimental validation are based.
@@
Moreover, despite recent efforts proposing various strategies to mitigate the generative catastrophic forgetting (GCF) in generative models, most remain misaligned with the core dynamics of diffusion-based architectures. Specifically, many methods rely on heuristics adapted from continual classification methods (Sun et al. 2024; Pfülb, Gepperth, and Bagus 2021; Varshney et al. 2021; Zhang et al. 2024a; Masip et al. 2023; Zając et al. 2023; Cheng et al. 2024), employ stability-plasticity trade-offs originally designed for Generative Adversarial Networks (GANs) (Ali, Rossi, and Bertozzi 2025; Ye and Bors 2021b, a; Gu et al. 2020; Zhao et al. 2024), or use trained diffusion models as ad hoc replay generators (Gao and Liu 2023). However, these techniques often conflict with the stochastic differential equations (SDEs) that govern the generative processes of diffusion models, leading to noticeable degradation in output quality. Empirical results confirm these limitations, as reflected in consistent declines in generative fidelity across sequential tasks, measured by the metric like Fréchet Inception Distance (FID) (Heusel et al. 2017) (see Figure. 1 and Table. CDG Pipeline). In response to these research limitations, we begin by analyzing the unique mechanisms of diffusion models and investigate how they can be optimized in the context of streaming tasks, aiming to preserve shared knowledge across tasks and mitigate GCF.
@@
To achieve this goal, we begin by formalizing diffusion trajectories, complex sequences of denoising operations that progressively transform noise into structured outputs, and examining how they interact through shared knowledge under streaming task scenarios. Grounded in Bayesian theory and multi-task learning principles (Yu et al. 2020), we derive a theoretical upper bound that quantifies the retention of generative knowledge across tasks in CDG, as formally defined in Theorem 1. Attaining this upper bound requires satisfying three critical consistency constraints: inter-task alignment of model-internalized knowledge, consistency in the mean embeddings of unconditional generated samples across tasks, and semantic consistency within the prior (i.e., label) space of original samples  [^2] Notably, the label space acts as a proxy for human prior knowledge, with the constraint aimed at preserving shared semantic structure across tasks, rather than improving classification performance. across tasks.
@@
Operationalizing the theoretical insights, we introduce the Continual Consistency Diffusion (CCD) framework, which translates the derived guarantees into a tractable, consistency-driven optimization strategy. CCD enforces cross-task stability through a hierarchical integration of consistency objectives. Specifically, Inter-task Knowledge Consistency regularizes model-internal knowledge representations across tasks, serving as the foundation for long-term retention. Unconditional Knowledge Consistency preserves intrinsic generative behavior in the absence of explicit human priors, ensuring temporal coherence in the denoising process. Complementarily, Prior Knowledge Consistency promotes alignment in the prior space by enforcing semantic correspondence between original samples across tasks. Collectively, these mechanisms move beyond standard regularization or classification heuristics. By directly constraining the geometric structure of the diffusion process, CCD enables robust and theoretically grounded continual generation in the CDG pipeline.
@@
In summary, our work bridges the critical gap between traditional static diffusion models and the dynamic nature of real-world data streams. We present three main contributions. First, to the best of our knowledge, we are the first to rigorously formulate the CDG pipeline from a diffusion perspective. Second, we establish the first theoretical framework for CDG rooted in SDE calculus, introducing novel stability bounds for SDE trajectories under sequential task adaptation (Theorem 1). Third, guided by these insights, we propose the CCD framework, which enforces intrinsic knowledge consistency through three synergistic components. Experiments on various benchmarks demonstrate its superiority, yielding significant gains while effectively mitigating GCF problem.
Related Works
@@
Diffusion Models. Diffusion-based generative architectures have redefined state-of-the-art performance in structured data synthesis, primarily due to their ability to invert stochastic denoising trajectories. At a foundational level, these models learn to reverse-engineer discrete Markov chains (DDPMs) (Ho, Jain, and Abbeel 2020) or continuous-time stochastic differential equations (SDEs) (Song et al. 2021). Recent advancements in noise scheduling (Nichol and Dhariwal 2021; Lu et al. 2022) and adaptive sampling techniques (Lu et al. 2022; Zheng et al. 2023) have further enhanced output fidelity. Such developments underscore diffusion models’ theoretical strength as universal data approximators (Song, Meng, and Ermon 2020), regardless of their discrete or continuous formulation (Bruce et al. 2024; Nie et al. 2025; Yang et al. 2025). However, their success heavily relies on closed-world assumptions, where training data remains static and entirely observable. Consequently, a critical challenge persists in adapting diffusion models to dynamic, streaming data environments, paralleling incremental human cognition.
Continual Classification. The continual classification (CC) aims to enable models to progressively acquire new classification knowledge while retaining previously learned information, addressing the challenge of Catastrophic Forgetting (CF) (Kirkpatrick et al. 2017; Li and Hoiem 2017; Parisi et al. 2019). Traditional CC methods include replay-based techniques, which store subsets of historical data to maintain stable performance (Rolnick et al. 2019), and regularization methods like EWC (Kirkpatrick et al. 2017) and LwF (Li and Hoiem 2017), which impose constraints on parameters to reduce interference between tasks. Additionally, gradient-based strategies such as GEM (Lopez-Paz and Ranzato 2017) orthogonalize gradients to minimize task conflicts. Recent advancements have focused on using generators trained on prior tasks as buffers, with DDGR (Gao and Liu 2023) as a prominent example. While effective in mitigating CF in CC tasks, DDGR incurs substantial training overhead due to the need to synthesize past samples during each training batch.
@@
In parallel, the advent of pre-training has driven progress in parameter-efficient fine-tuning techniques. Methods such as L2P (Wang et al. 2022b) and DualPrompt (Wang et al. 2022a) utilize task-specific prompts to effectively balance adaptability and knowledge retention. Techniques like S-Prompt (Wang, Huang, and Hong 2022) and CODA-Prompt (Smith et al. 2023) enhance performance by explicitly capturing domain relationships, while dynamic methods like DAP (Jung et al. 2023) and hierarchical approaches like HiDe-Prompt (Wang et al. 2024) support adaptation across diverse domains. Despite these advancements, the existing CC research remains focused on basic classification tasks, limiting applicability to complex real-world scenarios. To bridge this gap, we investigate extending CC methods to practical and challenging applications, namely continual generation, within our standardized CDG pipeline.
@@
Continual Generation. The continual generation represents a significant blind spot in the current landscape of CL research, with only a scant body of work dedicated to this area. Among these studies, most have primarily explored methods based on GAN architectures (Ali, Rossi, and Bertozzi 2025; Ye and Bors 2021b, a; Gu et al. 2020; Zhao et al. 2024). Additionally, many approaches attempt to adapt techniques originally developed for CC tasks (Sun et al. 2024; Pfülb, Gepperth, and Bagus 2021; Varshney et al. 2021; Zhang et al. 2024a; Masip et al. 2023), applying relevant fine-tuning strategies. However, given that generation and classification are fundamentally distinct tasks, such direct transfer is largely impractical. Our experimental findings further highlight that many methods effective in CC scenarios fail entirely in generation settings, sometimes even yielding adverse effects. This is primarily due to the substantial knowledge disparity across tasks, which causes the diffusion generator to collapse, a phenomenon we refer to as generative collapse. In this paper, we address this gap by developing a theoretical framework to model the task transition process of diffusion models.
Continual Consistency Diffusion
@@
In our standardized CDG pipeline, we consider a sequence of non-stationary tasks $\{\mathcal{T}_{k}\}_{k=1}^{K}$, where each task $\mathcal{T}_{k}$ has a distinct data distribution $p^{k}(x_{0})$ and corresponding label distribution $p^{k}(y|x_{0})$. The forward diffusion process for each task is governed by an SDE: $dx_{t}^{k}=f_{k}(x_{t}^{k},t)dt+g_{k}(t)dw_{t}$, where $x_{t}^{k}\in\mathbb{R}^{d}$ represents the diffused samples at time $t$ under task $\mathcal{T}_{k}$, and $g_{k}(t)$ controls time-dependent noise for each task. To build on the derivations in Appendix and enhance the preservation of shared knowledge across tasks, we employ a direct rehearsal buffer $\mathcal{B}^{real}_{k}=\{\hat{x}_{0},\hat{y}\sim p^{j}(x_{0},y)\}_{j=1}^{k-1}$,  [^3] We avoid generative replay as in DDGR (Gao and Liu 2023), since it requires synthesizing past samples for every training batch, leading to a 2–3× increase in optimization time and rendering it impractical despite its performance gains. which stores a limited set of real samples within a fixed storage budget $C$, following common practice in CL research such as (Li et al. 2025; Wan et al. 2025; Wang et al. 2025). This memory mechanism, both in its theoretical formulation and practical implementation, forms the core foundation for the retention of shared knowledge across tasks.
Building on these foundations, we now present a detailed exposition of our CCD optimization framework’s theory and loss formulation within the standardized CDG pipeline.
Theoretical Foundation
@@
A central challenge in CDG is formalizing the interaction of task-specific generative processes through shared SDE dynamics (Song et al. 2021), as in Equation 34 in the Appendix. Existing empirical approaches (Smith et al. 2024; Zhao et al. 2024; Zhang et al. 2024a) mainly adapt solutions from continual classification methods, but lack rigorous theoretical guarantees, often leading to GCF or rigid fixation. Our analysis begins by establishing fundamental bounds on cross-task knowledge retention, which are essential for systematic CDG. Through rigorous derivation, we establish Theorem 1, which lays the theoretical foundation for subsequent innovations.
Theorem 1 (Cross-Task Diffusion Evolution Bound)
@@
Let $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ represent two tasks in CDG, each characterized by distinct data distributions $p(x_{0})$ and $q(x_{0})$, along with their respective conditional prior distributions $p(y)$ and $q(y)$. The diffused processes for these tasks evolve over time as $\{p(x_{t})\}_{t=0}^{T}$ and $\{q(x_{t})\}_{t=0}^{T}$. Assume that for all $x_{0},y,t$, the conditional probability distributions of the two tasks satisfy $p_{t}(x_{t}|x_{0},y)=q_{t}(x_{t}|x_{0},y)$. Here, $p_{t}(x_{t}|x_{0},y)$ refers to the distribution $p(x_{t}|x_{0},y,t)$, and similarly, $q_{t}(x_{t}|x_{0},y)$ denotes $q(x_{t}|x_{0},y,t)$. Furthermore, let $\epsilon_{\theta}^{p}$ and $\epsilon_{\theta}^{q}$ represent the time-dependent score approximators, or noise estimators, for tasks $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. Under mild assumptions, we expect that the gradients of the mean functions $\mu(x_{t},t)$ and $\nu(x_{t},t)$ align, such that $\nabla_{x_{t}}\mu(x_{t},t)\approx\nabla_{x_{t}}\nu(x_{t},t)$, as noted in (Yu et al. 2020). Additionally, it is assumed that the variance $\sigma^{2}_{t}$ at any given time $t$ remains consistent across both tasks. Lastly, we assume that the evolving state $x_{t}$ does not influence the label $y$, meaning the label is independent of the diffusion process at any given time step. These conditions enable the potential retention and transfer of knowledge between tasks, leading to the derivation of an optimization upper bound for their interaction.  [^4] For the detailed proof, please see Appendix.
@@
$\mathcal{L}_{UB}\;=\;\kappa\mathcal{L}_{IKC}\;+\;\lambda\,\mathcal{L}_{UKC}\;+\;\eta\,\mathcal{L}_{PKC},$ (1)
@@
$\mathcal{L}_{IKC}\;=\;\epsilon_{\theta}^{q}(x_{t},y,t)\;-\;\epsilon_{\theta}^{p}(x_{t},y,t),$ (2)
@@
$\mathcal{L}_{PKC}\;=\;\frac{\bar{\alpha}_{t}}{\bar{\beta}_{t}}\mathbb{E}_{p_{t}(x_{0}|x_{t})}\Bigl{[}D_{\mathrm{KL}}\bigl{(}p_{t}(y|x_{0})\,\big{\|}\,q_{t}(y|x_{0})\bigr{)}\Bigr{]}.$ (4)
In particular, $\mathcal{L}_{UB}$ encapsulates three components: the inter-task knowledge consistency ($\mathcal{L}_{IKC}$), the unconditional knowledge consistency ($\mathcal{L}_{UKC}$), and the prior knowledge consistency ($\mathcal{L}_{PKC}$). Minimizing $\mathcal{L}_{UB}$ aligns the reverse-time diffusion gradients between tasks $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$, thereby allowing the two tasks to retain as much shared knowledge as possible during the SDE optimization process.
Basic Diffusion Model Training
To ensure the optimality of the score estimators $\epsilon_{\theta}^{p}$ and $\epsilon_{\theta}^{q}$ within $\mathcal{L}_{IKC}$ for effective subsequent retention and transfer, we define the fundamental training objective for each task $\mathcal{T}_{k}$. Following the standard DDPM framework (Ho, Jain, and Abbeel 2020; Song et al. 2021), the base objective for conditional generation minimizes the weighted $L_{2}$ error between predicted and actual noise:
@@
where $\bar{\alpha}_{t}$ and $\bar{\beta}_{t}$ follow the DDPM variance schedule (Ho, Jain, and Abbeel 2020), and $\epsilon$ is the standard Gaussian noise.
Building on this, to enhance the shared knowledge between different tasks, we incorporate the data from the rehearsal buffer into the training process. However, instead of random sampling, we concatenate the pairs $x_{0},\hat{x}_{0}>$, which also facilitates the implementation of Equations 2 to 4. Therefore, in CDG pipeline, the fundamental composite training objective for diffusion models can be expressed as: